OpenStudy (anonymous):

Use a system of equations to solve the word problem. A rectangular table has a perimeter of 18 feet. Its length is 5 feet greater than its width. Find the dimensions. A. l = 8 and w = 3 B. l = 7 and w = 2 C. l = 2 and w = 7 D. l = 3 and w = 8

OpenStudy (anonymous):

Well you can a automatically eliminate c. and d. because the length is not 5 feet longer than the width

OpenStudy (anonymous):

Then you just have to find which one works: 8 + 8 + 3 + 3 or 7 + 7 + 2 + 2

OpenStudy (anonymous):

So which one equals 18?

OpenStudy (anonymous):

7+7+2+2

OpenStudy (anonymous):

P = 2L + 2W L = W + 5 now sub W + 5 in for L 18 = 2(W + 5) + 2W 18 = 2W + 10 + 2W 18 = 4W + 10 18 - 10 = 4W 8 = 4W 2 = W L = W + 5 L = 2 + 5 L = 7 so width = 2 and length = 7